find measure of angle t if measure of angle t is 20 more than 4 times it supplement
Answer by John10(297)

The sum of two supplementary angles is 180 degrees.
t = 20 + 4(180 - t)
t = 20 + 720 - 4t
t + 4t = 740
5t = 740
t = 740/5
t = 148 degrees.
John10:)